OpenStudy (anonymous):

i forgot how to this this prom any help?? Compute: (4 + 9) − 3(6) ÷ 2 + 8

OpenStudy (anonymous):

You do what's inside the parentheses first. So add 4+9. Which gives us 13. Next work from left to right and do the multiplican and division next. So multiply 3(6) Which gives you 18. Next divide 18 by 2. 18/2=9 So your problem now looks like this: 13-9+8 Now work from left to right adding and subtracting the terms. So 13-9=4 This now gives us 4+8=12